integration of sin8x dx /(9+sin4x)1/2


 


plz explain in detail


Grade:12

1 Answers

Rahul askIITiansExpert.IITR
13 years ago
take sin2x=t

dt=2sinxcosxdx=sin2xdx

now the numerator of the expression: sin8xdx=2sin4xcos4xdx=4sin2xcos2xcos4xdx=4cos2xcos4xdt;

now we know that cos2x=cos2x-sin2x=1-2sin2x=1-2t

similarly, cos4x=1-sin22x=1-4sin2xcos2x=1-4t(1-t)=1-4t+4t2=4t2-2t-2t+1=(2t-1)2

the denominator is (9+t2)1/2

put the expressions in position

and expand the numerator

you will get different powers of t

separate than integrate each part separately

Cool

Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation.

DO APPROVE IF THE ANSWER HELPS YOU.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Regards,

Askiitians Experts

Rahul- IIT Roorkee